Write the equation of an ellipse, in standard form, with center at (-1,3), vertex at (3,3) and minor axis of length 2.
Answer by Boreal(15235)

The vertex is 4 units from the center and the minor axis is 2 units long, so the semi-minor axis is 1 unit.
a^2 will be 16 and b^2=1
{(x+1)^2/16} + {(b-3)^2)/1}=1
a^2-c^2=b^2
16-c^2=1
c^2=15
c= +/- sqrt (15). Those are foci.
{(x+1)^2/16} + {(b-3)^2)/1}=1